# Lifting the Veil on the Large Language Model Supply Chain: Composition, Risks, and Mitigations

### v2

@@ 4.3.3. Risks on Models
Model Stealing. Model stealing refers to attacks where an adversary attempts to extract sensitive information about a machine learning model (e.g., model gradients, training data, and model parameters), thereby compromising the model’s privacy. It encompasses various model privacies, including gradient leakage (Gupta et al., 2022), training data extraction (Zhang et al., 2023b; Yang et al., 2023; Yu et al., 2023), and model information extraction (Carlini et al., 2024). It not only infringes the intellectual property rights of the model owner but also leads to security risks. For example, adversaries can create shadow models via model stealing, which are then used to generate adversarial examples, reverse-engineer sensitive data used in model training (Pan et al., 2020), or bypass any access restrictions or usage limitations in the original model (Shen et al., 2023). It can cause the privacy risk of model stealing (PR3).
Regarding LLMs, several model-stealing attacks have been reported to be effective (Zhang et al., 2023b; Yang et al., 2023; Parikh et al., 2022). Li et al. and Yu et al. (Carlini et al., 2021; Yu et al., 2023) conducted experiments to investigate the effectiveness of training data extraction on LLMs. Truong et al. (Truong et al., 2021) proposed techniques from the domain of data-free knowledge transfer to perform model extraction, successfully replicating the capabilities of the original model. Carlini et al. (Carlini et al., 2024) demonstrated precise and significant information extraction from black-box generative language models by recovering the embedding projection layer of a transformer model, given typical API access.
4.3.4. Risks on Prompts
Since LLMs are typically built from general knowledge, prompts are an effective way to incorporate domain-specific insights, directly enhancing the quality, relevance, and accuracy of the model’s output. The shared prompts allow a wider community to contribute and share specialized knowledge. However, shared prompts present several risks. In the context of shared prompts, those jailbreak prompts can be maliciously leveraged and pose challenges to the security mechanism of existing LLMs.
Implant Poisoned Prompts. Implanting poisoned prompts deliberately inserting harmful or misleading prompts into publicly accessible repositories or datasets. These poisoned prompts can be designed to manipulate the behavior of large language models (LLMs) by influencing their output in undesirable ways. Since shared prompts are widely shared and reused, such poisoned prompts can easily propagate through the supply chain, leading to inaccurate, biased, or even harmful model responses for downstream users who incorporate them into their applications. It can cause the performance degradation (DR1) of dilivery risks.
  • Prompt injection. Prompt injection is a type of attack that exploits malicious prompt words with trusted prompts against LLM applications. By exploiting the incapability of distinguishing “good” and “bad” prompts, prompt injection can cause malicious inputs to affect the model’s behavior and output, leading to information leakage or other security vulnerabilities. Stumpp et al.(Greshake et al., 2023) demonstrated the potential of prompt injection to indirectly execute arbitrary code within a Python interpreter in LLM applications, thereby gaining control over the host server (e.g., accessing environment variables, arbitrary file contents, executing DoS attacks, etc.). Chen et al.(Chen et al., 2024) and Shayegani et al. (Shayegani et al., 2023) further explored the threats of prompt injection using multi-modal inputs.
  • Jailbreaking. Jailbreaking attempts to subvert safety filters built into the LLMs themselves (Willison, 2024). Recent research (Wei et al., 2024; Deng et al., 2024; Zou et al., 2023) has used a variety of methods to make the LLMs LLMs responsive to restricted or insecure content (e.g., content that contains pornography, biased content, or assists the user in committing criminal acts). Wei et al. (Wei et al., 2024) investigates the reasons for the success of such attacks and how they occur, emphasizing the need for safety-capability parity. To assess the potential harm caused by jailbreaking, Shen et al. (Shen et al., 2023) devised the JailbreakHub framework including 1,405 jailbreak prompts.
Figure 6. Risky Actions, Supply Chain Components and Risk Types Introduced by Risky Consumers
4.3.5. Risks on LLM Applications
The LLM application performs various tasks using the LLM involving natural language processing, task-based dialogue, database search, etc. It can be a web application, an agent or a chatbot. The potential risks may be exposed in multiple ways.
Exploit LLM Application Vulnerabilities. The LLM Plugin Store (e.g., GPT store (OpenAI, 2024)) is a platform where users can discover, install, and manage plugins designed for enhancing the functionality of large language models. However, attackers can publish malicious plugins, guiding the model to output incorrect content through prompt injection (Wu et al., 2024). Moreover, some third-Party plugins could lead to account takeovers and data leakage (Newsroom, 2024; Kovacs, 2024a, b). Exploit lLM application vulnerabilities can potentially cause the hacker use (SR1), data leakage (PR1) and performance degradation (DR1).
Excessive Access. Users may overload the LLM application with excessive requests, potentially causing it to malfunction or become unresponsive. A Denial of Service (DoS) attack is a type of cyberattack that seeks to deplete computational resources, resulting in latency or rendering resources unavailable. Given that large language models (LLMs) demand substantial computational power, attackers can deliberately craft prompts to diminish the availability of these models (Derner et al., 2023). Shumailov et al. (Shumailov et al., 2021) demonstrated the feasibility of performing DoS attacks within the domain of LLMs. These attacks significantly increased the energy consumption and latency of the model (boosted to 10 to 200 times), raising researchers’ concerns about performance security in critical decision-making scenarios. Stumpp et al. (Stumpp, 2024) et al. used prompt injection to control the execution of dead-end code by the code executor of an LLM application, resulting in an unresponsive program host server. Excessive access can potentially cause hacker use (SR1), system compromise (SR2), emergence of abnormal content (SR3), data leakage (PR1), personally identifiable information (PII) leakage (PR2), model stealing (PR3), code leakage (PR4), and prompt leakage (PR5), and performance degradation (DR1).
Feedback Pollution. Existing LLM systems are often tweaked based on user feedback to improve their performance (Griffith et al., 2013). However, attackers may disrupt the performance of an LLM system by injecting a large amount of irrelevant data, leading to a surge in false feedback. Consequently, analysts at the victim organization may find themselves spending excessive time reviewing and rectifying erroneous inferences (MITRE, 2021). More seriously, LLM model developers may use this erroneous feedback during the instruction tuning phase to conduct reinforcement learning based on human feedback. This situation can not only lead to a decline in model performance but also cause the model to exhibit unpredictable behavior when facing real-world data (Wan et al., 2023; Baumgärtner et al., 2024). Sun et al. (Sun et al., 2024) demonstrate two proof-of-concept examples, stealthily manipulating a neural code generation system to generate code with vulnerabilities, attack payload, and malicious and spam messages. Feedback pollution can potentially cause the hacker use (SR1), system compromise (SR2), emergence of abnormal content (SR3), privacy risk (PR).
@@ 4.3.5. Risks on LLM Applications
Perform Middleman Attack. A Man-in-the-Middle (MitM) attack involves an attacker intercepting communication between two parties without their knowledge. In the context of using the LLM Application or any chatbot platform, a MitM attack could occur if an attacker gains access to the communication channel, allowing them to alter or steal sensitive information being shared between a user and the AI system. Kang et al. (Kang et al., 2024) demonstrated that attackers may intercept data transmitted during a user’s interaction with LLM-based applications, leading to unauthorized access to sensitive information. Si et al. (Si et al., 2023) proposed that attackers can set up a fake API that mimics a legitimate one while offering a lower price to lure users. It not only allows the attacker to modify or steal sensitive data but also to exploit economic gains from the middle. The middleman attack can potentially cause data leakage (PR1), personally identifiable information (PII) leakage (PR2), model stealing (PR3), code leakage (PR4), and prompt leakage (PR5).
@@ 4.3.6. Risks on Platforms
Many platforms play important roles in the LLM supply chain, supporting various stages such as source code management, data hosting, model training, monitoring, and deployment. However, these platforms can introduce inherent risks.
@@
Camouflage Contributors. Attackers can create fake accounts that imitate existing platform accounts (e.g., official accounts) to publish malicious data, code, models, and more. Users may be misled by these repositories, inadvertently using malicious tools to develop or deploy large models. It can cause the risk of malicious use (SR1).
@@
Malicious Contributors. Attackers can obtain and exploit credentials of existing accounts or API tokens through various methods (News, 2018, 2022). An LLM red team organization (Team, 2024) has demonstrated in practice that once in possession of these credentials, attackers can gain unauthorized access to a range of resources and services associated with LLMs, potentially compromising the integrity, confidentiality, and availability of the targeted systems and data. Attackers can also use social engineering tactics to become maintainers of legitimate projects. Once they have obtained maintainer status, they can introduce malicious code or dependencies, which may go unnoticed by other contributors or users due to the perceived legitimacy of the new maintainer. For instance, in the XZ attack, the attacker gained maintainer access to the XZ project, embedding a backdoor in a critical dependency (Bals, 2024). It can cause the risk of malicious use (SR1).
Exploit Platform Vulnerabilities. Exploiting platform vulnerabilities refers to attackers identifying and taking advantage of weaknesses in the underlying infrastructure of an LLM or related application. These vulnerabilities can exist in the platform’s software, hardware, or configuration, allowing attackers to perform unauthorized actions such as gaining access to sensitive data, altering model behavior, or disrupting services. It can cause the risk of hacker use (SR1), data leakage (PR1) and model stealing (PR3).
  • CI/CD Flaws. Attackers can leverage flaws in AI cloud services to carry out attacks. Researchers from Wiz identified vulnerabilities in SAP’s AI Core service, which could have allowed attackers to access sensitive data of other tenants within SAP’s cloud infrastructure (scmagazine, 2024). By exploiting these flaws, attackers can retrieve AWS tokens, access AI training data, and potentially cause data theft, service disruption, and even supply chain attacks within LLM applications.
  • Privacy Leakage. Attackers can exploit secrets found in public artifacts generated by automated CI/CD pipelines. Recent research revealed that artifacts produced by GitHub Actions workflows may contain sensitive tokens, such as GITHUB_TOKEN and ACTIONS_RUNTIME_TOKEN, which attackers could use to inject malicious code (French, 2024). When users download and utilize these compromised artifacts, it can result in a supply chain attack.
Figure 7. Risky Actions, Supply Chain Components and Risk Types Introduced by Risky Users
4.3.7. General Risks to LLM Supply Chain Component
@@
Apart from risks that are specific to the six composing elements, we elaborate on general risks to LLM supply chain components.
@@
Introduce Risky Components. A benign contributor may not realize that the components they introduce have hidden vulnerabilities or security flaws. These risky components can be exploited by malicious actors once deployed. For example, an LLM application might rely on a library that appears secure but contains a hidden vulnerability. When the benign contributor integrates this library, they inadvertently expose the LLM and its users to potential exploitation. Attackers can leverage these risky components for data theft, model manipulation, or unauthorized access. It can cause the risk of hacker use (SR1), data leakage (PR1) and model stealing (PR3).
@@
Discontinue Maintenance. When a contributor stops maintaining a component, updates for security patches, bug fixes, and compatibility improvements are halted. It can lead to vulnerabilities if the abandoned component contains outdated dependencies that are prone to exploitation. Moreover, the discontinuation can disrupt further development by creating compatibility issues, reducing model performance, or complicating integration with newer systems. Furthermore, it poses significant risks because malicious actors could take advantage of this by forking the abandoned project and inserting malicious code. It can cause the risk of unmaintained (DR2).
@@
Release Vulnerable Code. The benign contributors may submit code that contains security flaws or exploits without malicious intent. Such vulnerabilities could range from weak encryption practices, insecure data handling, or exploitable logic errors. While the contributor’s intent is not harmful, the inadvertent introduction of security issues can still have serious consequences, potentially leading to supply chain attacks. It can cause hacker use (SR1), system compromise (SR2), and the emergence of abnormal content (SR3).
@@
Raise License Disputes. While open-source licenses generally grant permission to use, modify, and distribute the code, the owners retain the copyright. It introduces potential risks for users who download and use the code, model or data if contributors claim copyrights on their code. They may enforce restrictions or take legal action if users violate the terms of the license. In 2020, CoKinetic Systems sued Panasonic for allegedly violating the GPLv2 license, with the lawsuit seeking financial compensation (Murray, 2020). It can cause the risk of license disputes (LR1).
Figure 8. Risky Actions, Supply Chain Components and Risk Types Introduced by Risky Administrators
5. Mitigation of Risks in Large Language Model Supply Chain
We elaborate on the mitigation of risks in the LLM supply chain with regard to the seven aspects.
5.1. Mitigation to Risks from Third-party Libraries
@@
Malware Detection. Malware detection identifies viruses, worms, trojans, and backdoors from third-party libraries in both binaries and source code. Key techniques include signature comparison, static analysis, and dynamic analysis. Signature-based detection tools, like VirusTotal (VirusTotal, 2023), are effective for identifying binary malware. To detect malware within package registries, static and dynamic analyses are widely applied. Recent research has advanced these approaches by modeling the behavior of malicious packages using static characteristics alone (Huang et al., 2024b; Zhang et al., 2023a), as well as integrating both static and dynamic analyses to enhance detection accuracy (Duan et al., 2020; Huang et al., 2024a). This measure can mitigate the risk SR2.
Software Bill-of-Materials Analysis. A Software Bill of Materials (SBOM) is a list of inventory components included in imported third-party libraries. SBOM analysis provides visibility into all third-party libraries and their versions, helping users avoid unvetted or potentially risky libraries. Several commercial tools offer SBOM analysis, including Sonatype Management (Sonatype, 2008), BlackDuck (BlackDuck, 2002), and Microsoft’s SBOM tool (Microsoft, 2022). By utilizing SBOM analysis, users can enhance security and improve overall software quality throughout the development lifecycle. This measure can mitigate the risk SR2 and SR3.
5.2. Mitigation to Risks from Dataset
@@
Data-Deduplication. In the face of privacy risks, we can leverage data deduplication on the dataset. Jagielski et al. discovered that model privacy attacks are more effective on the training dataset that appears multiple times (Jagielski et al., 2023). Removing duplicate text from the training data  (Abadi et al., 2024) helps protect against model memorization of sensitive information, making privacy attacks less effective. This measure can mitigate the risk PR2.
@@
Data Sanitization. Data sanitization cleans input data to defend against data poisoning and backdoor attacks by applying techniques like data denoising, filtering, and smoothing, which effectively remove adversarial noise while retaining legitimate information  (Xu et al., 2017). Additionally, these methods help prevent data leakage by ensuring that shared or processed data cannot be traced back to individuals. This measure can mitigate the risk SR3 and PR1.
Data Watermark. Data watermark protects data privacy by embedding specific watermark identifiers (Sun et al., 2023; Tang et al., 2023) into the training dataset. During the training process, the model is trained using watermarked data, ensuring that it generates output text containing the corresponding watermark when given specific inputs, thus achieving the goal of identifying and tracing the data source. This measure can mitigate the risk PR1.
@@
5.3. Mitigation to Model Risks

### v3

@@ 5.4.1. Risks on Artifact
Steal Models. Adversaries extract sensitive information (e.g., model gradients, model parameters) of machine learning models (Gupta et al., 2022; Zhang et al., 2023b; Yang et al., 2023; Yu et al., 2023; Carlini et al., 2024b), which are then used to generate adversarial examples and bypass any access restrictions (Shen et al., 2023). It can cause privacy risks (PR3).
5.4.2. Risks on LLM Apps
The LLM app can be a web application, an agent or a chatbot.
@@ 5.4.2. Risks on LLM Apps
Perform Middleman Attacks. A Man-in-the-Middle (MitM) attack involves an attacker intercepting communication between two parties without their knowledge. MitM attack in the LLM application occurs if an attacker gains access to the communication channel, allowing them to alter or steal sensitive information being shared between a user and the AI system (Kang et al., 2024; Si et al., 2023). It can cause data leakage (PR1), personally identifiable information (PII) leakage (PR2), model stealing (PR3), code leakage (PR4), and prompt leakage (PR5).
Exploit LLM Application Vulnerabilities. The LLM Plugin Store (e.g., GPT store (OpenAI, 2024)) enables users to install plugins to enhance large language model functionalities. However, attackers can publish malicious plugins, guiding the model to output incorrect content through prompt injection (Wu et al., 2024), lead to account takeovers and data leakage (Newsroom, 2024; Kovacs, 2024a, b). It can potentially cause the hacker use (SR1), data leakage (PR1) and performance degradation (DR1).
Perform Excessive Access. The LLM application may be unresponsive or malfunction due to excessive requests. For example, performing denial of service attacks within the domain of LLMs can significantly increase the energy consumption and latency of the model (Shumailov et al., 2021). It can cause hacker use (SR1), system compromise (SR2), emergence of abnormal content (SR3), data leakage (PR1), PII leakage (PR2), model stealing (PR3), code leakage (PR4), prompt leakage (PR5), and performance degradation (DR1).
Perform Feedback Pollution. LLM applications typically leverage user feedbacks to improve their performance (Griffith et al., 2013) (e.g., reinforcement learning based on human feedback), which can possibly be disrupted by false feedbacks (Wan et al., 2023; Baumgärtner et al., 2024). It can potentially cause the hacker use (SR1), system compromise (SR2), emergence of abnormal content (SR3), privacy risk (PR).
@@ 5.4.3. Risks on Platforms
The platforms support various usages for LLM application such as source code management, data hosting, model training, monitoring, and deployment.
@@
Camouflage Contributors. Attackers can create fake accounts that imitate existing platform accounts (e.g.,, official accounts) to publish malicious data, code, models, and more. Users may be misled by these repositories, inadvertently using malicious tools to develop or deploy large models. It can cause the risk of malicious use (SR1).
@@
Introduce Malicious Contributors. Attackers can obtain and exploit credentials of existing accounts or API tokens through various methods (News, 2018, 2022) (e.g., social engineering). Once in possession of these credentials, attackers can gain unauthorized access to a range of resources and services associated with LLMs (Team, 2024). It can cause the risk of malicious use (SR1).
Exploit Platform Vulnerabilities. Attackers can identify and take advantage of weaknesses in the underlying platforms. They leverage flaws in cloud services (e.g., SAP AI service (scmagazine, 2024)) to exploit cross-session information leakage, or retrieve sensitive tokens found in CI/CD artifacts (e.g., GitHub Actions (French, 2024)). It can cause the risk of hacker use (SR1), data leakage (PR1) and model stealing (PR3).
5.4.4. Risks on Toolchain
Malicious risks can be introduced at both the model development and deployment phases, compromising the integrity and security of the entire pipeline.
Implant Malicious Code. Attackers can compromise the toolchain by injecting malicious code into the toolchain. For instance, ”Xcode-Ghost” (wikipedia, 2024), a malicious version of Xcode, was distributed through unofficial channels. Developers unknowingly used this compromised version to build their apps, resulting in the inclusion of malicious code in the final applications. It can cause the security risk of system compromise (SR2) and the privacy risk (PR).
Exploit Tool vulnerabilities. Attackers exploit known vulnerabilities in development tools, frameworks, or CI/CD components that have not been promptly patched. For instance, the vulnerability of Jenkins Core (jenkins, 2024) can lead to remote code execution (RCE) via agent connections. It can cause the security risk of system compromise (SR2), Performance Degradation(DR1) and Denial of Service (DR3) by triggering unexpected errors or system failures.
5.4.5. Generic Risks to LLM Components
@@
Apart from risks that are specific to the four types of components, we elaborate on generic risks to LLM components.
@@
Introduce Risky Components. A naive benign contributor may not be fully aware of the hidden vulnerabilities or security flaws in LLM components. When the benign contributor integrates this component, they inadvertently expose the LLM and its users to potential exploitation. It can cause the risk of hacker use (SR1), data leakage (PR1) and model stealing (PR3).
@@
Discontinue Maintenance. When a contributor stops maintaining a component, updates for security patches, bug fixes, and compatibility improvements are halted. The abandoned component contains outdated and incompatible dependencies that are prone to exploitation or reduce model performance and reliability. It can cause the risk of unmaintained (DR2).
@@
Release Vulnerable Code. The benign contributors may submit code that contains security flaws or exploits without malicious intent, ranging from weak encryption practices, insecure data handling, or exploitable logic errors. It can cause hacker use (SR1), system compromise (SR2), and the emergence of abnormal content (SR3).
@@
Raise License Disputes. While open-source licenses grant general permissions, the owners retain the copyright of the artifact. It introduces potential risks for users who download and use the code, model or data if contributors claim copyrights on their code. They may enforce restrictions or take legal action if users violate the terms of the license (Murray, 2020). It can cause the risk of license disputes (LR1).
Summary. We illustrate the risky scenarios, including high-risk stakeholders, actions, risk types, and their corresponding components within the supply chain. Compared to the traditional software supply chain, the LLM supply chain involves unique risks, such as implanting poisoned data, model stealing, and model inference attacks. Some risk types are exclusive to LLMs, such as model leakage. The affected components in the LLM supply chain represent a new range of software elements that deserve further investigation and consolidation.
6. Mitigation of Risks in Large Language Model Supply Chain (RQ3)
We elaborate on the mitigation of risks in the LLM supply chain with regard to six aspects. Figure 9 presents the distribution of mitigation strategies discussed in the surveyed literature. The majority of studies propose solutions based on proactive detection (74 papers, 33.2%) and protective operation (56 papers, 25.1%). Resilient training techniques also account for a significant portion, reflecting efforts to enhance model robustness during development. The fourth and fifth most discussed strategies involve management (20 papers, 9.0%) practices and accountability mechanisms (17, 7.6%), which primarily address socio-organizational aspects of risk mitigation. Notably, only one paper discusses the use of contingency planning as a mitigation approach.
Figure 8. Distribution of Mitigations that Covered in our Collected Literature
Figure 9. Mitigations against Risky Actions and Risk Types
6.1. Proactive Detection
@@
Malware Detection. Malware detection identifies viruses, worms, trojans, and software backdoors from third-party libraries in both binaries and source code. Key techniques include signature comparison, static analysis, and dynamic analysis. Signature-based detection tools, like VirusTotal (VirusTotal, 2023), are effective for identifying binary malware. Recent research has advanced these approaches by modeling the behavior of malicious packages using static characteristics alone (Huang et al., 2024b; Zhang et al., 2023a), as well as integrating both static and dynamic analyses to enhance detection accuracy (Duan et al., 2020; Huang et al., 2024a). This measure can mitigate the risk SR2.
Sanitization. Sanitization involves various techniques to enhance data and prompt security, mitigating risks SR3, PR1, and PR2.
  • Data Sanitization. Data sanitization cleans input data by applying techniques like data denoising, filtering, and smoothing, which effectively remove adversarial noise while retaining legitimate information  (Xu et al., 2017). These methods also help prevent data leakage by ensuring that shared or processed data cannot be traced back to individuals.
@@
  • Data Deduplication. Jagielski et al. discovered that model privacy attacks are more effective on the training dataset that appears multiple times (Jagielski et al., 2023). Removing duplicate text from the training data  (Abadi et al., 2024) helps protect against model memorization of sensitive information.
  • Prompt Sanitization. Prompt sanitization filters, formats, or regularized prompts are provided to LLMs to ensure that the model does not misinterpret their original intent, thereby preventing abnormal content or malicious behavior. Suo et al. (Suo, 2024) applied to sign to sensitive elements (e.g.,’delete file’) in the user’s prompt and Robey et al.  (Robey et al., 2023) randomly perturb characters in a given input prompt (e.g., inserting random characters), effectively prevent jailbreaking in LLMs. Chen et al .(Chen et al., 2024) separates prompts and user data into two independent channels to prevent prompt injection.
6.2. Protective Operation
@@
Access Control. Access control mechanisms are essential measures to ensure systems secure interactions and protect sensitive information in systems.
  • Account Access Control. Authorization (i.e., restricting access to authorized users), authentication (i.e., enforcing mutual authentication and encryption), and rate Limiting (i.e., limiting access to verified users via rate-limiting mechanisms) ensure authorized clients, credible feedback and controlled access rate to prevent system abuse (e.g., Denial of Service (DoS) attacks). This measure can mitigate the risks DR1, DR2 and DR3.
  • Session and Data Isolation. Session isolation techniques (e.g., sandboxing) separates user sessions, preventing cross-session attacks. Besides, platforms can utilize secrets management tools, log scanning to ensure that sensitive information like tokens and API keys are not exposed in public repositories or logs (French, 2024). This measure can mitigate the risks SR2, PR1, PR2, PR3 and PR4.
Homomorphic Encryption. Homomorphic encryption is a form of encryption that allows computations to be performed directly on encrypted data without needing to decrypt it first. Chen et al (Chen et al., 2022) employs homomorphic encryption to enable privacy-preserving inference for Transformer-based models. This measure can mitigate the risk PR1.
Multi-Party Computation. Li et al. (Li et al., 2022) achieve private transformer inference using Secure Multi-Party Computation. Federated learning can be regarded as a particular type of multi-party computation. It enables multiple clients collaboratively train a shared model without transferring their individual data to a central server (Li et al., 2020; Zhang et al., 2022) while keeping the shared model remain secure and resistant to backdoor attacks or malicious data manipulation. This measure can mitigate the risk PR1, PR3 and SR3.
Vulnerability Detection. Vulnerability detection identifies weaknesses or flaws in software or systems that could be exploited by attackers. For example, platforms should also ensure that their systems are following best security practices, such as least privilege, secure credential storage, and network segmentation, keeping all CI/CD tools and dependencies up to date with the latest security patches to protect against known vulnerabilities (scmagazine, 2024). This measure can mitigate the risk SR2.
6.3. Accountability Mechanism
Software Bill-of-Materials Analysis. A Software Bill of Materials (SBOM) is a list of inventory components included in imported third-party libraries, helping users avoid unvetted or potentially risky libraries. Several commercial tools offer SBOM analysis, including Sonatype Management (Sonatype, 2008), BlackDuck (BlackDuck, 2002), and Microsoft’s SBOM tool (Microsoft, 2022). Given the widespread use of large language model applications, SBOM analysis can be expanded to encompass not only third-party libraries, but also models, data, prompts, and other related elements. This measure can mitigate the risk SR2 and SR3.
Watermark. Watermark techniques embed identifiable information into data or models to ensure traceability, authenticity, and protection against misuse.
  • Data Watermark. Data watermark protects data privacy by embedding specific watermark identifiers (Sun et al., 2023; Tang et al., 2023) into the training dataset, thus identifying and tracing the data source. This measure can mitigate the risk PR1.
  • Model Watermark. Model watermark embeds traceable information into the text generated by LLMs to verify the model’s source and authenticity without degrading output quality, such as logits modification (Kirchenbauer et al., 2023), entropy-based embedding (Lee et al., 2023), multi-bit payload (Wang et al., 2023), and token sampling (Kuditipudi et al., 2023). This measure can mitigate the risk PR3.